OpenStudy (anonymous):

expand the logarithm using the three laws of logarithms. log2 (ab squared)

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Rule number 1:\[\Large\rm \log_{2}\left(\color{royalblue}{x}\color{orangered}{y}\right)=\log_{2}\left(\color{royalblue}{x}\right)+\log_{2}\left(\color{orangered}{y}\right)\]Do you understand how we can apply that rule to our problem?\[\Large\rm \log_{2}\left(\color{royalblue}{a}\color{orangered}{b^2}\right)=?\]